Sheffield Gravity Grips produces spike sets for track shoes. CEO Brittany Sheffield has gathered the following information about the company’s sales volume and marketing cost for the past six months:

Sales Volume Total Marketing Costs
January 550,928 $82,744
February 390,732 $76,214
March 561,372 $83,040
April 543,000 $82,489
May 546,656 $82,488
June 553,288 $82,908

Using the high-low method, compute the variable marketing cost per spike set. (Round unit cost to 2 decimal places, e.g.52.75.)

A. Variable Marketing Costs = $____ Per spike set sold.

B. Compute the total fixed marketing costs.

C. Represent the Marketing Cost Function in equation form.

D. Examine the data and identify the potential outlier.

E. Recalculate the marketing cost function, removing the potential outlier.
Answer #1
Marketing Costs
Sales volume Marketing Costs
High level of activity(March)                               5,61,372 $83,040
Low level of activity(February)                               3,90,732 $76,214
Change                               1,70,640 $6,826
Variable cost per unit =$6,826 / 170,640 units =$0.04 per unit
Fixed cost element =$83,040 - (561,372*$0.04) =$60,585
The equation is:
Y =$60,585 + $0.04X
In the given data the Sales and costs for February can be identified as an outlier.
So noe the revised cost function can be found as follows:
Marketing Costs
Sales volume Marketing Costs
High level of activity                               5,61,372 $83,040
Low level of activity                               5,43,000 $82,489
Change                                  18,372 $551
Variable cost per unit =$551 / 18,372 =$0.03 per unit
Fixed cost element =$83,040 - (561,372*$0.03) =$66,199
The equation is:
Y =$66,199 + $0.03 or 66,200 + $0.03X